OpenStudy (hannaha):

MEDAL FOR HELP!! PLEASE HELP EXPLAIN THIS!! AS IN HOW TO DO IT--NOT JUST THE ANSWER!! Simplify this expression. Type the variable term before the constant. 18 + 5(2x - 1) + 3

OpenStudy (hannaha):

Was just about to tag you @rockycatsrule

OpenStudy (rockycatsrule):

lol

OpenStudy (rockycatsrule):

r u soving for x or for y

OpenStudy (hannaha):

I'm simplifying

OpenStudy (hannaha):

So not trying to solve for anything

OpenStudy (rockycatsrule):

Let's simplify step-by-step. 18+5(2x−1)+3 Distribute: =18+(5)(2x)+(5)(−1)+3 =18+10x+−5+3 Combine Like Terms: =18+10x+−5+3 =(10x)+(18+−5+3) =10x+16 Answer: =10x+16

OpenStudy (hannaha):

Just combining like terms

OpenStudy (rockycatsrule):

yep

OpenStudy (anonymous):

Do the multiplication first: 5(2x-1) = 10x-5 Then add up the integer numbers, 10, 18, -5

OpenStudy (rockycatsrule):

does that some it up

OpenStudy (hannaha):

so 10x + 23?

OpenStudy (rockycatsrule):

yep

OpenStudy (hannaha):

Wait bc your answer said 10x + 16?

OpenStudy (anonymous):

10x + 16

OpenStudy (hannaha):

ok but 10 + 18 +-5 is 23

OpenStudy (anonymous):

18 + 10x -5 +3

OpenStudy (rockycatsrule):

i hop that i was good help and u now understand it a little better

OpenStudy (hannaha):

oh so you're not adding the 10 sorry duh

OpenStudy (rockycatsrule):

lol yea

OpenStudy (hannaha):

Thanks you guys

OpenStudy (rockycatsrule):

welcome